# Model Predictive Control via Probabilistic Inference: A Tutorial and Survey

### v2

@@ 3.1.3 Guiding Samples with Auxiliary Controllers
When linear approximations of dynamics and cost are available, analytic feedback controllers offer a computationally cheap solution. Examples include task-specific Linear Quadratic Regulators (LQR) [trevisan2024biased], iterative Linear Quadratic Gaussian (iLQG) controllers [williams2018robust, gandhi2021robust], and covariance steering controllers [yin2022trajectory]. For problems where such approximations are unavailable, gradient-based methods provide a more general alternative: convex optimization [rastgar2024priest], Stochastic Gradient Descent (SGD) [bharadhwaj2020model], and Adam [heetmeyer2023rpgd] can guide samples using only gradient information of the cost and dynamics. For the most general setting, i.e., where even cost and dynamics gradients may not be available, Stein Variational Gradient Descent (SVGD) offers a derivative-free option that approximates the KL gradient from samples alone [lambert2021stein, honda2024stein, miura2024spline].
@@ 3.1.4 Cost-Aware Adaptive Prior Distribution
The central difficulty is a circular dependency: evaluating the quality of a prior requires drawing i.i.d. samples and rolling them out to compute trajectory costs, but these costs are themselves influenced by the prior. Asmar et al.  [asmar2023model] break this circularity through iterative refinement, i.e., alternating between cost evaluation and prior update, at the expense of higher computational cost. Lighter-weight alternatives avoid full iteration: Tao et al.  [tao2022path] use Stochastic Control Barrier Functions (SCBF) to set a safe mean and variance, while Honda et al.  [honda2024stein] employ SVGD on a small set of pilot samples to capture the mode shape before the main sampling phase. Other approaches sidestep the circularity by predicting future cost characteristics: Mohamed et al.  [mohamed2025towards] propagate variance dynamics via the unscented transform and fold the result into the cost, and Yi et al.  [yi2024covo] derive a closed-form optimal covariance under simplifying assumptions on the cost structure.
@@ 3.1.5 Learning-based Prior Distribution
When offline data or simulation is available, machine learning can be used to construct priors that are already well-aligned with the task before online execution begins.
If expert demonstrations are available, generative models trained on the dataset can serve directly as priors. Urain et al.  [urain2022learning] learn an energy-based model of the optimal trajectory distribution, while Carvalho et al.  [carvalho2023motion] use diffusion models [ho2020denoising] to capture richer multi-modal structure. A complementary line of work incorporates model-free RL policies as priors: pre-trained policies can initialize the MPPI mean [qu2023rl, wang2025residualmppi], or MPPI can be embedded as a residual correction during RL training so that the learned policy and the MPC prior co-evolve [sacks2024deep]. In model-based RL, MPPI serves as the planner while a model-free policy provides the prior [hansen2022temporal].
More specialized approaches learn environment-specific priors using Normalizing Flows [power2022variational, sacks2023learning]. By incorporating the KL objective in Eq. (4) into the training loss, these methods can learn priors in a self-supervised manner without expert data, conditioning only on environment features and start–goal pairs.
@@ 3.2.1 Variational Inference with Multi-modal Variational Distributions
The most direct strategy is to replace the unimodal Gaussian of MPPI with a more expressive variational family. Osa et al.  [osa2020multimodal] formulated offline trajectory optimization as cost-weighted density estimation using Gaussian Mixture Models (GMMs), and Okada et al.  [okada2020variational] adopted GMMs as the variational distribution within the PI-MPC framework, optimizing via mirror gradient descent. An alternative to parametric mixtures is to represent the variational distribution nonparametrically as a set of particles updated via SVGD [lambert2021stein, barcelos2021dual]. Wang et al.  [wang2021variational] generalized these particle-based methods by replacing the KL divergence with the Tsallis divergence, providing additional flexibility in controlling mode-covering behavior. Outside the strict VI framework, Carvalho et al.  [carvalho2023motion] integrated diffusion models into PI-MPC to generate multi-modal trajectory samples.
The primary strength of these methods is their ability to maintain solution diversity, which enhances exploration and can improve learning efficiency in model-based RL [okada2020variational]. Their main weakness relative to MPPI is computational: as shown in Eq. (12), MPPI obtains the globally optimal Gaussian approximation through a single round of sampling and weighted averaging, whereas multi-modal VI methods require iterative parameter updates, increasing both computational cost and the risk of numerical instability [^7] MPPI also requires iterative updates when simultaneously optimizing the mean and covariance [wang2021variational]..
@@ 3.2.2 Seeking a Single Mode in Multi-modal Distributions
Clustering-based methods achieve this by partitioning samples post hoc. Park et al.  [park2025csc] employed DBSCAN to cluster samples by density, computed a weighted average within each cluster, and selected the cluster representative with the lowest cost. Zhang et al.  [zhang2024multi] computed separate weighted averages for multiple cost components and then re-weighted all samples to fuse them into a single control sequence.
A more principled alternative is to change the divergence objective. Standard PI-MPC minimizes the forward KL divergence $D_{\mathrm{KL}}(\pi^{*}\|\pi_{\boldsymbol{\theta}})$, which encourages the variational distribution to cover all modes (mode-covering). Minimizing the reverse KL divergence $D_{\mathrm{KL}}(\pi_{\boldsymbol{\theta}}\|\pi^{*})$ instead yields a mode-seeking solution that concentrates on a single mode. Because the reverse KL does not admit a closed-form solution, iterative methods such as SVGD [honda2024stein] and mirror descent [kobayashi2022real] are used in practice.
@@ 3.3 Handling of Constraints
Real-world optimal control problems invariably involve constraints on both states and control inputs. Control input constraints (e.g., actuator limits) are relatively easy to enforce: samples that exceed bounds can be clipped after drawing from the prior [williams2018information], or inputs can be mapped through saturating functions such as $\tanh$; either approach can be absorbed into the dynamics of Eq. (1b). State constraints (e.g., collision avoidance, joint limits) are fundamentally harder because they couple the constraint with the forward dynamics and cannot be enforced by modifying a single sample in isolation. The simplest remedy is to add penalty terms to the cost function in Eq. (1a)—PI-MPC readily accommodates discontinuous penalties such as indicator functions and cost maps [macenski2020marathon2]—but penalties cannot guarantee strict satisfaction: if all samples from the prior violate a constraint, the “optimal” solution will itself be infeasible. Methods for stricter enforcement fall into two categories: (i) projecting samples onto the feasible set during prior sampling, and (ii) incorporating constraints directly into the variational inference problem. Projection methods are simpler to implement but may require differentiability or linearity assumptions; constrained VI is more principled but computationally heavier.
@@ 3.3.1 State-constrained Prior Sampling
Several projection mechanisms have been proposed. Control Barrier Functions (CBFs) [gandhi2023safe, gandhi2022safety] project inputs to enforce forward-invariance of a safe set; primal–dual interior-point methods [park2025csc] and convex optimization [rastgar2024priest] solve small local programs to find the nearest feasible input. Yan et al.  [yan2024output] take the inverse approach: they sample directly in the state space and convert feasible state trajectories to control inputs via inverse dynamics.
For settings that require long-horizon safety guarantees, Borquez et al.  [borquez2025dualguard] project samples onto a future safety region defined by Hamilton–Jacobi reachability analysis, ensuring that the system can always return to a safe state. Wang et al.  [wang2025constrained] extend the framework to hierarchical equality and inequality constraints by combining the Lagrange multiplier method with null-space projection.
@@ 3.3.2 Constrained Particle-based Variational Inference
A more principled alternative is to embed state constraints directly into the variational inference problem, yielding a constrained KL minimization over Eq. (9). This line of work builds on constrained VI methods from the machine learning literature [liu2021sampling, zhang2022sampling], often implemented via SVGD [liu2016stein], which represents the target distribution as particles and updates them along the gradient of the KL divergence.
Power et al.  [power2024constrained] extended the constrained SVGD framework of [zhang2022sampling] to handle multiple linear and nonlinear constraints, applying it to robotic planning problems by projecting particle updates onto the tangent space of the constraint manifold. Tabor et al.  [tabor2025constrained] proposed a more unified formulation that folds constraints into the SVGD kernel, avoiding explicit tangent-space projection.
@@ 3.4 Applications to High-Dimensional Robots
Early PI-MPC studies focused on low-dimensional systems such as ground vehicles with simple action spaces. As the field moves toward more capable platforms. e.g., humanoids with 30+ joints, multi-fingered hands, and soft robots, the dimensionality of the control problem grows rapidly, and naive Monte Carlo sampling becomes prohibitively expensive. Scaling PI-MPC to high degrees of freedom (DoF) requires addressing two coupled challenges: the exponential growth of sample complexity with action-space dimension, and the difficulty of obtaining accurate dynamics models for systems with complex contact and deformation. This section reviews how recent work tackles these challenges for platforms including 4-Wheel Independent Drive and Steering (4WIDS) vehicles [aoki2024switching], manipulators [bhardwaj2022storm], quadruped robots [keshavarz2025control], and humanoid robots [xue2024full].
@@ 3.4.1 Sample Complexity in High DoF Robots
The most common countermeasure is dimensionality reduction. Defining the action space in terms of spline control points, rather than per-timestep joint commands, reduces the effective dimensionality and has enabled PI-MPC on various humanoid platforms [howell2022predictive, pezzato2025sampling, miura2024spline]. Combining Halton sequences (for low-discrepancy coverage) with spline interpolation further improves sample efficiency [bhardwaj2022storm, pezzato2025sampling], though such smoothing techniques inevitably sacrifice some optimality by restricting the space of representable trajectories.
An alternative to dimensionality reduction is iterative refinement. Diffusion-style annealing [pan2024model] and adaptive importance sampling (AIS) [asmar2023model] sharpen the sample distribution over multiple passes, concentrating samples in high-quality regions. These methods have shown promising results, including real-world deployment on quadruped [keshavarz2025control] and humanoid [xue2024full] robots, but the additional iterations increase computational cost.
@@ 3.4.2 Dynamics Modeling for High DoF Robots
Accurate dynamics models are essential for PI-MPC, yet analytical modeling becomes intractable for high-DoF robots, especially those involving contact (e.g., legged locomotion, manipulation). A practical solution is to use GPU-accelerated physics simulators, e.g., MuJoCo [todorov2012mujoco], Isaac Lab [mittal2025isaaclab], and Brax [freeman1brax], for forward rollouts during planning [howell2022predictive, pezzato2025sampling, xue2024full]. Isaac Lab and Brax are particularly well-suited to PI-MPC because they support massively parallel simulation on GPUs, and their randomizable environments enable domain randomization for robustness.
An alternative is to learn the dynamics model from data and use it for planning, i.e., the model-based RL. Methods such as TD-MPC [hansen2022temporal, hansentd] learn compact latent-space dynamics and plan with MPPI, achieving strong performance even on high-DoF continuous-control tasks.
@@ 3.5 Hardware Acceleration
Real-time PI-MPC demands that all computation, i.e., sampling, rollout, cost evaluation, and weight update, complete within a few milliseconds. Because the sampling and rollout steps are embarrassingly parallel, the primary lever for meeting this budget is hardware-level parallelism. The choice of platform depends on the deployment context: CPUs suffice for low-dimensional systems with moderate sample counts, GPUs are the current workhorse for high-dimensional problems, and emerging architectures (e.g., FPGAs, quantum-inspired hardware) target ultra-low-latency or energy-constrained settings.
On the CPU side, the ROS 2 Navigation stack includes an MPPI-based local planner [macenski2020marathon2] implemented with the Eigen library, offering efficient single-threaded and SIMD-parallel execution suitable for embedded deployment. On the GPU side, machine-learning frameworks such as PyTorch and JAX provide high-level abstractions that make GPU-accelerated MPPI straightforward to implement [bou2023torchrl], though Python-level overhead limits throughput. Native CUDA kernel implementations in C++ eliminate this overhead and deliver substantially higher performance [vlahov2024mppi].
Beyond conventional CPUs and GPUs, recent work has begun exploring unconventional accelerators. Werthen-Brabants and Simoens et al.  [werthen2025ising] proposed solving the MPPI optimization on Ising machines, a form of quantum-inspired computing. Tanguy-Legac et al.  [tanguy2026domain] investigated dedicated FPGA and ASIC designs tailored to the MPPI dataflow.
@@ 3.6 Theoretical Analysis of MPPI
Despite its empirical success, PI-MPC lacks the mature theoretical toolkit available for gradient-based MPC or LQR. We must often tune parameters by trial and error, without formal guidance on how close the resulting solution is to the true optimum or how many samples are sufficient. The two central questions are: (i) how large is the optimality gap between the MPPI solution and the solution of the original optimal control problem, and (ii) under what conditions does MPPI converge, and at what rate? Progress on these questions not only informs algorithm design but is also essential for certifying PI-MPC in safety-critical applications.
@@ 3.6.1 Optimality Gap from Original Optimal Control Problem
Because MPPI minimizes the KL objective in Eq. (5) rather than the original cost in Eq. (1), its solution can deviate from the true optimum. Williams et al.  [williams2018information] identified a special case in which this gap vanishes: when the cost separates into state-dependent and input-dependent terms and the input cost is a quadratic weighted by the prior covariance, the MPPI solution coincides exactly with the optimal solution of the original problem. Trevisan et al.  [trevisan2024biased] generalized this result, showing that for any variational family, adding an appropriate bias correction to the cost function eliminates the gap.
More broadly, the optimality gap can be traced to the prior regularization term in Eq. (5). Appropriate prior design (Section 3.1) and temperature tuning can therefore reduce the gap in practice. Yi et al.  [yi2024covo] derived a closed-form covariance that maximizes the convergence rate, directly linking prior design to optimality. Homburger et al.  [homburger2025optimality] provided a more general analysis for deterministic problems, proving that as $\lambda\to 0$ the MPPI solution converges to the deterministic optimum and bounding the control-input and value-function errors as functions of $\lambda$.

### v3

@@ 3.1.3 Guiding Samples with Auxiliary Controllers
When linear approximations of dynamics and cost are available, analytic feedback controllers offer a computationally cheap solution. Examples include task-specific Linear Quadratic Regulators (trevisan2024biased), iterative Linear Quadratic Gaussian (iLQG) controllers (williams2018robust; gandhi2021robust), and covariance steering controllers (yin2022trajectory). For problems where such approximations are unavailable, gradient-based methods provide a more general alternative: convex optimization (rastgar2024priest), Stochastic Gradient Descent (bharadhwaj2020model), and Adam (heetmeyer2023rpgd) can guide samples using only gradient information of the cost and dynamics. For the most general setting, i.e., where even cost and dynamics gradients may not be available, Stein Variational Gradient Descent (SVGD) offers a derivative-free option that approximates the KL gradient from samples alone (lambert2021stein; honda2024stein; miura2024spline).
@@ 3.1.4 Cost-Aware Adaptive Prior Distribution
The central difficulty is a circular dependency: evaluating the quality of a prior requires drawing i.i.d. samples and rolling them out to compute trajectory costs, but these costs are themselves influenced by the prior. asmar2023model break this circularity through iterative refinement, i.e., alternating between cost evaluation and prior update, at the expense of higher computational cost. Lighter-weight alternatives avoid full iteration: tao2022path use Stochastic Control Barrier Functions (SCBF) to set a safe mean and variance, while honda2024stein employ SVGD on a small set of pilot samples to capture the mode shape before the main sampling phase. Other approaches sidestep the circularity by predicting future cost characteristics: mohamed2025towards propagate variance dynamics via the unscented transform and fold the result into the cost, and yi2024covo derive a closed-form optimal covariance under simplifying assumptions on the cost structure.
@@ 3.1.5 Learning-based Prior Distribution
When offline data or simulation is available, machine learning can be used to construct priors that are already well-aligned with the task before online execution begins. If expert demonstrations are available, generative models trained on the dataset can serve directly as priors. urain2022learning learn an energy-based model of the optimal trajectory distribution, while carvalho2023motion use diffusion models (ho2020denoising) to capture richer multi-modal structure. A complementary line of work incorporates model-free RL policies as priors: pre-trained policies can initialize the MPPI solution (qu2023rl; wang2025residualmppi), or MPPI can be embedded as a residual correction during RL training so that the learned policy and the MPC prior co-evolve (sacks2024deep). In model-based RL, MPPI serves as the planner while a model-free policy provides the prior (hansen2022temporal).
More specialized approaches learn environment-specific priors using Normalizing Flows (power2022variational; sacks2023learning). By incorporating the KL objective in Eq. (4) into the training loss, these methods can learn priors in a self-supervised manner without expert data, conditioning only on environment features and start–goal pairs.
@@ 3.2.1 Variational Inference with Multi-modal Variational Distributions
The most direct strategy is to replace the unimodal Gaussian of MPPI with a more expressive variational family. osa2020multimodal formulated offline trajectory optimization as cost-weighted density estimation using Gaussian Mixture Models (GMMs), and okada2020variational adopted GMMs as the variational distribution within the PI-MPC framework, optimizing via mirror gradient descent. An alternative to parametric mixtures is to represent the variational distribution nonparametrically as a set of particles updated via SVGD (lambert2021stein; barcelos2021dual). wang2021variational generalized these particle-based methods by replacing the KL divergence with the Tsallis divergence, providing additional flexibility in controlling mode-covering behavior. Outside the strict VI framework, carvalho2023motion integrated diffusion models into PI-MPC to generate multi-modal trajectory samples.
The primary strength of these methods is their ability to maintain solution diversity, which enhances exploration and can improve learning efficiency in model-based RL (okada2020variational). Their main weakness relative to MPPI is computational: as shown in Eq. (12), MPPI obtains the globally optimal Gaussian approximation through a single round of sampling and weighted averaging, whereas multi-modal VI methods require iterative parameter updates, increasing both computational cost and the risk of numerical instability [^7] MPPI also requires iterative updates when simultaneously optimizing the mean and covariance (wang2021variational)..
@@ 3.2.2 Seeking a Single Mode in Multi-modal Distributions
Clustering-based methods achieve this by partitioning samples post hoc. park2025csc employed DBSCAN to cluster samples by density, computed a weighted average within each cluster, and selected the cluster representative with the lowest cost. zhang2024multi computed separate weighted averages for multiple cost components and then re-weighted all samples to fuse them into a single control sequence.
A more principled alternative is to change the divergence objective. Standard PI-MPC minimizes the forward KL divergence $D_{\mathrm{KL}}(\pi^{*}\|\pi_{\boldsymbol{\theta}})$, which encourages the variational distribution to cover all modes (mode-covering). Minimizing the reverse KL divergence $D_{\mathrm{KL}}(\pi_{\boldsymbol{\theta}}\|\pi^{*})$ instead yields a mode-seeking solution that concentrates on a single mode. Because the reverse KL does not admit a closed-form solution, iterative methods such as SVGD (honda2024stein) and mirror descent (kobayashi2022real) are used in practice.
@@ 3.3 Handling of Constraints
Real-world optimal control problems invariably involve constraints on both states and control inputs. Control input constraints (e.g., actuator limits) are relatively easy to enforce: samples that exceed bounds can be clipped after drawing from the prior (williams2018information), or inputs can be mapped through saturating functions such as the hyperbolic tangent function; either approach can be absorbed into the dynamics of Eq. (1b). State constraints (e.g., collision avoidance, joint limits) are fundamentally harder because they couple the constraint with the forward dynamics and cannot be enforced by modifying a single sample in isolation. The simplest remedy is to add penalty terms to the cost function in Eq. (1a)—PI-MPC readily accommodates discontinuous penalties such as indicator functions and cost maps (macenski2020marathon2)—but penalties cannot guarantee strict satisfaction: if all samples from the prior violate a constraint, the “optimal” solution will itself be infeasible. Methods for stricter enforcement fall into two categories: (i) projecting samples onto the feasible set during prior sampling, and (ii) incorporating constraints directly into the variational inference problem. Projection methods are simpler to implement but may require differentiability or linearity assumptions; constrained VI is more principled but computationally heavier.
@@ 3.3.1 State-constrained Prior Sampling
Several projection mechanisms have been proposed. Control Barrier Functions (gandhi2023safe; gandhi2022safety) project inputs to enforce forward-invariance of a safe set; primal–dual interior-point methods (park2025csc) and convex optimization (rastgar2024priest) solve small local programs to find the nearest feasible input. yan2024output take the inverse approach: they sample directly in the state space and convert feasible state trajectories to control inputs via inverse dynamics.
For settings that require future safety guarantees, borquez2025dualguard project samples onto a future safety region defined by Hamilton–Jacobi reachability analysis, ensuring that the system can always return to a safe state. wang2025constrained extend the framework to hierarchical equality and inequality constraints by combining the Lagrange multiplier method with null-space projection.
@@ 3.3.2 Constrained Particle-based Variational Inference
A more principled alternative is to embed state constraints directly into the variational inference problem, yielding a constrained KL minimization over Eq. (9). This line of work builds on constrained VI methods from the machine learning literature (liu2021sampling; zhang2022sampling), often implemented via SVGD (liu2016stein), which represents the target distribution as particles and updates them along the gradient of the KL divergence.
power2024constrained extended the constrained SVGD framework of zhang2022sampling to handle multiple linear and nonlinear constraints, applying it to robotic planning problems by projecting particle updates onto the tangent space of the constraint manifold. tabor2025constrained proposed a more unified formulation that folds constraints into the SVGD kernel, avoiding explicit tangent-space projection.
@@ 3.4 Applications to High-Dimensional Robots
Early PI-MPC studies focused on low-dimensional systems such as ground vehicles with simple action spaces. As the field moves toward more capable platforms. e.g., humanoids with 30+ joints, multi-fingered hands, and soft robots, the dimensionality of the control problem grows rapidly, and naive Monte Carlo sampling becomes prohibitively expensive. Scaling PI-MPC to high degrees of freedom (DoF) requires addressing two coupled challenges: the exponential growth of sample complexity with action-space dimension, and the difficulty of obtaining accurate dynamics models for systems with complex contact and deformation. This section reviews how recent work tackles these challenges for platforms including 4-Wheel Independent Drive and Steering (4WIDS) vehicles (aoki2024switching), manipulators (bhardwaj2022storm), quadruped robots (keshavarz2025control), and humanoid robots (xue2024full).
@@ 3.4.1 Sample Complexity in High DoF Robots
The most common countermeasure is dimensionality reduction. Defining the action space in terms of spline control points, rather than per-timestep joint commands, reduces the effective dimensionality and has enabled PI-MPC on various humanoid platforms (howell2022predictive; pezzato2025sampling; miura2024spline). Combining Halton sequences (for low-discrepancy coverage) with spline interpolation further improves sample efficiency (bhardwaj2022storm; pezzato2025sampling), though such smoothing techniques inevitably sacrifice some optimality by restricting the space of representable trajectories.
An alternative to dimensionality reduction is iterative refinement. Diffusion-style annealing (pan2024model) and adaptive importance sampling (AIS) (asmar2023model) sharpen the sample distribution over multiple passes, concentrating samples in high-quality regions. These methods have shown promising results, including real-world deployment on quadruped (keshavarz2025control) and humanoid (xue2024full) robots, but the additional iterations increase computational cost.
@@ 3.4.2 Dynamics Modeling for High DoF Robots
Accurate dynamics models are essential for PI-MPC, yet analytical modeling becomes intractable for high-DoF robots, especially those involving contact (e.g., legged locomotion, manipulation). A practical solution is to use GPU-accelerated physics simulators, e.g., MuJoCo (todorov2012mujoco), Isaac Lab (mittal2025isaaclab), and Brax (freeman1brax), for forward rollouts during planning (howell2022predictive; pezzato2025sampling; xue2024full). Isaac Lab and Brax are particularly well-suited to PI-MPC because they support massively parallel simulation on GPUs, and their randomizable environments enable domain randomization for robustness.
An alternative is to learn the dynamics model from data and use it for planning, i.e., the model-based RL. Methods such as TD-MPC (hansen2022temporal; hansentd) learn compact latent-space dynamics and plan with MPPI, achieving strong performance even on high-DoF continuous-control tasks.
@@ 3.5 Hardware Acceleration
Real-time PI-MPC demands that all computation, i.e., sampling, rollout, cost evaluation, and weight update, be completed within a few milliseconds. Because the sampling and rollout steps are embarrassingly parallel, the primary lever for meeting this budget is hardware-level parallelism. The choice of platform depends on the deployment context: CPUs suffice for low-dimensional systems with moderate sample counts, GPUs are the current workhorse for high-dimensional problems, and emerging architectures (e.g., FPGAs, quantum-inspired hardware) target ultra-low-latency or energy-constrained settings.
On the CPU side, the ROS 2 Navigation stack includes an MPPI-based local planner (macenski2020marathon2) implemented with the Eigen library, offering efficient single-threaded and SIMD-parallel execution suitable for embedded deployment. On the GPU side, machine-learning frameworks such as PyTorch and JAX provide high-level abstractions that make GPU-accelerated MPPI straightforward to implement (bou2023torchrl), though Python-level overhead limits throughput. Native CUDA kernel implementations in C++ eliminate this overhead and deliver substantially higher performance (vlahov2024mppi).
Beyond conventional CPUs and GPUs, recent work has begun exploring unconventional accelerators. werthen2025ising proposed solving the MPPI optimization on Ising machines, a form of quantum-inspired computing. tanguy2026domain investigated dedicated FPGA and ASIC designs tailored to the MPPI dataflow.
@@ 3.6 Theoretical Analysis of MPPI
Despite its empirical success, PI-MPC lacks the mature theoretical toolkit available for gradient-based MPC. We must often tune parameters by trial and error, without formal guidance on how close the resulting solution is to the true optimum or how many samples are sufficient. The two central questions are: (i) how large is the optimality gap between the MPPI solution and the solution of the original optimal control problem, and (ii) under what conditions does MPPI converge, and at what rate? Progress on these questions not only informs algorithm design but is also essential for certifying PI-MPC in safety-critical applications.
@@ 3.6.1 Optimality Gap from Original Optimal Control Problem
Because MPPI minimizes the KL objective in Eq. (5) rather than the original cost in Eq. (1), its solution can deviate from the true optimum. williams2018information identified a special case in which this gap vanishes: when the cost separates into state-dependent and input-dependent terms and the input cost is a quadratic weighted by the prior covariance, the MPPI solution coincides exactly with the optimal solution of the original problem. trevisan2024biased generalized this result, showing that for any variational family, adding an appropriate bias correction to the cost function eliminates the gap.
More broadly, the optimality gap can be traced to the prior regularization term in Eq. (5). Appropriate prior design (Section 3.1) and temperature tuning can therefore reduce the gap in practice. yi2024covo derived a closed-form covariance that maximizes the convergence rate, directly linking prior design to optimality. homburger2025optimality provided a more general analysis for deterministic problems, proving that as $\lambda\to 0$ the MPPI solution converges to the deterministic optimum and bounding the control-input and value-function errors as functions of $\lambda$.